OpenStudy (anonymous):

how do i solve -xsquared=-5x+6

OpenStudy (anonymous):

add x squared to both sides of the equation and post what you got.

OpenStudy (anonymous):

0=xsquared-5x+6

OpenStudy (anonymous):

now, do you know how to solve quadratic equations of the form ax2+bx+c = 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then go ahead and do that in your problem.

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

mary, we help them get their answers, not solve their problems for them

OpenStudy (anonymous):

a=0xsquared b=-5x and c=6

OpenStudy (anonymous):

connor that is not the answer so keep working

OpenStudy (anonymous):

how did you get a = 0?

OpenStudy (anonymous):

-xsquared=-5x+6 +x +x -5x+6+x -6 -6 -5x-5 /-5 /-5 x=1

OpenStudy (anonymous):

that it

OpenStudy (anonymous):

0 = x^2 - 5x + 6 that factors into 0 = (x - 3)(x - 2) do you know how to solve from here?

OpenStudy (anonymous):

sorry internet stopped working

OpenStudy (anonymous):

so -2,-1?

OpenStudy (anonymous):

either x - 3 = 0 or x - 2 = 0 solve for x x = 3 or x = 2

OpenStudy (anonymous):

but im supposed to get the coordinates like (blank,0) and (blank,0)

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

so -2,0 and -3,0?